Solve the linear system:
y=2x+1
-2x+3y=-9

Answer by scott8148(6628)

substituting the value of y from the first equation into the second __ -2x+3(2x+1)=-9

-2x+6x+3=-9 __ 4x=-12 __ x=-3

y=2(-3)+1 __ y=-5